OpenStudy (anonymous):

okay what the heck am i supposed to do here! cos beta= -.82; 180 degrees< beta<270 degrees

OpenStudy (anonymous):

I is supposed to solve the equation how do I do that?

hartnn (hartnn):

even u getting beta value as 145.08 degrees from calculator?

OpenStudy (anonymous):

umm please give me a second

OpenStudy (anonymous):

no i got 0.999897589371505 but i just typed it in normal how do you do beta value?

hartnn (hartnn):

cos ^-1 of -0.82 if u get in radians, convert in degree

OpenStudy (anonymous):

why cos-1?

hartnn (hartnn):

no u should do cos inverse as cos beta =-0.82 beta = arccos (-0.82) or cos^-1 (-0.82)

OpenStudy (anonymous):

i'm still not that good with inverseing why do we do that for?

hartnn (hartnn):

if \(sin x=\theta\) then \(\theta = sin^{-1}x\)

hartnn (hartnn):

to get the angle.

OpenStudy (anonymous):

please hold i need to re check my calc

OpenStudy (anonymous):

okay so now i soo how you got 145.08 degrees what is it for?

hartnn (hartnn):

i used calculator what u got when u used calculator?

OpenStudy (anonymous):

same i just want to know what we do with it now cuz it's supposed to be > 180 and < 270

hartnn (hartnn):

okk, we use this cos(360-theta)=cos theta put theta=145.08 here

OpenStudy (anonymous):

ok so 360-145.08=214.92 is that it?

hartnn (hartnn):

yup thats your required angle.

OpenStudy (anonymous):

so just to be clear we subtract theta from 360 because 360 is like the circle and theta is like the piece missiing from it so we get the degree of wha'ts left as our answer yes?

hartnn (hartnn):

u can say that..... but be careful with sine sin(360-x)=- sin x sin(-x)=-sin x cos(-x)=cos x cos(360-x)=cos x

hartnn (hartnn):

as 0 degree corresponds to 360 degree

OpenStudy (anonymous):

so does tan(360-x)=tan x?

hartnn (hartnn):

its -tan x because tan x = sin x/cos x

OpenStudy (anonymous):

tan(360+x)=tan x tan(-x)=-tan x So tan(360-x)=tan(-x+360)=tan(-x)=-tan x

OpenStudy (anonymous):

wait i think i remember something like this so tan=sin/cos and you inverse sin and cos right?

hartnn (hartnn):

inverse sin and cos? why?to do what?

OpenStudy (anonymous):

no like if it say sin beta=x then you do sin^-1x and ummm.... idk bout cos

hartnn (hartnn):

yes, for any ratio u can take inverse if tan beta= x beta = arctan x if cos beta=y beta = arccos y

OpenStudy (anonymous):

awesomeness i think i get it but lets check

OpenStudy (anonymous):

tan beta= -5.3;90 degrees < beta < 180 so i do tan^-1(-5.3) yes?

hartnn (hartnn):

yup,do it

OpenStudy (anonymous):

okay please hold

OpenStudy (anonymous):

i got –79.31 so now i just subtract that from 360 and..... okay i got 439.31 and I know that can't be right... help?

hartnn (hartnn):

but u can't use 360 with tan because it becomes negative tan(360-x)=-tan x here u use 180 as tan's periodicity is 180 tan(180+x)=tan x now put x=-79.31 and tell me what u get

OpenStudy (anonymous):

i got 100.69!

hartnn (hartnn):

that is so correct :)

hartnn (hartnn):

if u want to verify tan 100.69=-5.3

OpenStudy (anonymous):

so i apologize but why is tan 180 instead of 360 because the other two are

hartnn (hartnn):

because of the periodicity sine and cos have 360 degrees as periodicity means values for sin and cos repeat after 360 sin(x)=sin(x+360) cos(x)=cos(x+360) whereas tan has periodicity of 180 all tan values repeat after 180 degrees tan(180+x)= tan x

OpenStudy (anonymous):

oh does this has anything to do with the unit circle?

hartnn (hartnn):

yup,thats related http://en.wikipedia.org/wiki/File:Unit_circle_angles_color.svg u can see graph of tan here: http://en.wikipedia.org/wiki/File:Tan(theta).jpg

OpenStudy (anonymous):

cool i got it thank you and have a good night ^-^

hartnn (hartnn):

welcome :) i'll have good evening here.